# ”Security for Everyone” in Finite Blocklength IRS-aided Systems With Perfect and Imperfect CSI

### v2

@@ III-1 Sub-Problem for Optimizing the Beamforming Vectors
Finally, by substitution (22) in (16), and introducing an auxiliary variable $\Gamma$ as the lower bound of the FBR-SR, we can recast problem ($\mathcal{P}1.1$) as:
$(\mathcal{P}1.3):$ $\max_{\Gamma,\boldsymbol{w}},$ (23a)
s.t. $\Gamma\leq\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta}^{(\iota)}),\leavevmode\nobreak\ \forall k,$ (23b)
($,$ (23c)
@@
problem ($\mathcal{P}1.3$) is an SDP problem with its linearized objective function in (22). This problem can be solved using standard solvers, e.g., the interior point method or the CVX toolbox [30].
@@ III-2 Sub-Problem for Optimizing the PREs
Likewise, given $\boldsymbol{w}$, we aim to find $\boldsymbol{\theta}^{(\iota+1)}$ such that, $\mathcal{S}_{k}(\boldsymbol{w}^{(\iota+1)}_{k},\boldsymbol{\theta}^{(\iota+1)})\geq\mathcal{S}_{k}(\boldsymbol{w}^{(\iota+1)}_{k},\boldsymbol{\theta}^{(\iota)}).$ Similar to the previous section, the lower bound approximation of the user’s SR can be obtained using the inequality (74) in Appendix D. The user’s data rate can be expressed as in (24), where
${q}_{1,k}^{(\iota+1)}$ $\triangleq C_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})-\gamma_{k}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})-\sigma_{k}{n}_{1,k}^{(\iota+1)},$
${n}_{1,k}^{(\iota+1)}$ $\triangleq{1}/{\rho_{k}^{(\iota+1)}}-{1}/{\upsilon_{{}_{k}}^{(\iota+1)}},$
${\boldsymbol{m}}^{(\iota+1)}_{k,k}(n)$ $\triangleq{\hat{\boldsymbol{m}}^{(\iota+1)}_{k,k}(n)}/{\rho_{k}^{(\iota+1)}},$
$\hat{\boldsymbol{m}}^{(\iota+1)}_{k,k}(n)$ $\triangleq\left(\boldsymbol{w}^{(\iota+1)}_{k}\right)^{H}\boldsymbol{h}_{{}_{k}}^{H}(\boldsymbol{\theta}^{(\iota)}){\boldsymbol{u}}_{k}\boldsymbol{\Xi}_{n}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}^{(\iota+1)}_{k},$
$\boldsymbol{\varphi}_{1,k}^{(\iota+1)}$ $\triangleq-{n}_{1,k}^{(\iota+1)}{\sum}_{j=1}^{\mathcal{K}}\boldsymbol{\varphi}_{k,j}^{(\iota+1)},$
$\boldsymbol{\varphi}_{k,j}^{(\iota+1)}$ $\triangleq\left(\boldsymbol{\aleph}_{k,j}^{(\iota+1)}(n)\right)^{*}\boldsymbol{\aleph}_{k,j}^{(\iota+1)}(m),n\in N,m\in N,$
$\boldsymbol{\aleph}_{k,j}^{(\iota+1)}(n)$ $\triangleq\boldsymbol{u}_{k}\boldsymbol{\Xi}_{n}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}^{(\iota+1)}_{j}.$
@@
$q_{1,k_{e}}^{(\iota+1)}$ $\triangleq\ln\left(\rho_{e}^{(\iota+1)}\right)-\rho_{e}^{(\iota+1)}-\ln\left(\upsilon_{e}^{(\iota+1)}\right)+1,$
$n_{1,k_{e}}^{(\iota+1)}$ $\triangleq\left(-\frac{\rho_{e}^{(\iota+1)}}{1+\rho_{e}^{(\iota+1)}}-\frac{1}{\upsilon_{e}^{(\iota+1)}}\right)\sigma_{e},$
${\boldsymbol{m}}^{(\iota+1)}_{k_{e}}(n)$ $\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}{\boldsymbol{m}}^{(\iota+1)}_{j,k_{e}}(n),$
${\boldsymbol{m}}^{(\iota+1)}_{j,k_{e}}(n)$ $\triangleq\left(\boldsymbol{w}^{(\iota+1)}_{k}\right)^{H}\boldsymbol{h}_{e}^{H}(\boldsymbol{\theta}^{(\iota)})\boldsymbol{u}_{e}\boldsymbol{\Xi}_{n}{\boldsymbol{G}}_{{}_{\text{{AR}}}}\boldsymbol{w}^{(\iota+1)}_{k},$
$\boldsymbol{\varphi}_{1,k_{e}}^{(\iota+1)}$ $\triangleq\left(\frac{-\rho_{e}^{(\iota+1)}}{(1+\rho_{e}^{(\iota+1)})}-\frac{1}{(\sigma_{e}+\upsilon_{e}^{(\iota+1)})}\right)\left(\sum\limits_{j=1}^{\mathcal{K}}{\boldsymbol{\varphi}}_{k_{e},j}^{(\iota+1)}\right),$
$\boldsymbol{\varphi}_{k_{e},j}^{(\iota+1)}$ $\triangleq\left({\boldsymbol{\aleph}}_{k_{e},j}^{(\iota+1)}(n)\right)^{*}\boldsymbol{\aleph}_{k_{e},j}^{(\iota+1)}(m),$
$\boldsymbol{\aleph}_{k_{e},j}^{(\iota+1)}(n)$ $\triangleq\boldsymbol{u}_{e}\boldsymbol{\Xi}_{n}{\boldsymbol{G}}_{{}_{\text{{AR}}}}\boldsymbol{w}^{(\iota+1)}_{j}.$
Next, we express the user-$k$’s and Eve dispersion factor. For $i\in\{k,e\}$, we can express the dispersion factors as in (26), where $q_{i,k}^{(\iota+1)}$ is expressed as in (27), $\boldsymbol{d}_{i,j}^{(\iota+1)}(n)$ is expressed as in (28), and
$\boldsymbol{d}_{i}^{(\iota+1)}$ $\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}\boldsymbol{d}_{i,j}^{(\iota+1)},$
$\boldsymbol{\varphi}_{2,i}^{(\iota+1)}$ $\triangleq{n}_{2,i}^{(\iota+1)}{\sum}_{j=1}^{\mathcal{K}}\boldsymbol{\varphi}_{i,j}^{(\iota+1)},$
${n}_{2,i}^{(\iota+1)}$ $\triangleq{(\xi_{i}\rho^{(\iota+1)}_{i})}/{\left(\left(\upsilon^{(\iota+1)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)})}\right)}.$
@@
$\mathcal{S}_{k}^{\mathcal{F}}\left(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}\right)$ $\geq q_{k}^{(\iota+1)}+2\sum\limits_{n=1}^{N}\Re\left\{\boldsymbol{m}^{(\iota+1)}_{k}(n)e^{j\theta_{n}}\right\},$ (29)
where $q_{k}^{(\iota+1)}$ is expressed as in (30), ${\boldsymbol{m}}^{(\iota+1)}_{k}(n)$ is expressed as in (31), and
$\boldsymbol{\varphi}_{k}^{(\iota+1)}$ $\triangleq\boldsymbol{\varphi}_{1,k}^{(\iota+1)}+\boldsymbol{\varphi}_{1,k_{e}}^{(\iota+1)}+\boldsymbol{\varphi}_{k_{e},j}^{(\iota+1)}+\boldsymbol{\varphi}_{2,k}^{(\iota+1)}+\boldsymbol{\varphi}_{2,k_{e}}^{(\iota+1)}.$
With (29), we formulate the following optimization problem to obtain $\boldsymbol{\theta}^{(\iota+1)}$,
$(\mathcal{P}1.4):$ $\underset{\boldsymbol{\theta}}{\max}\leavevmode\nobreak\ \underset{k\in\mathcal{K}}{\min}\leavevmode\nobreak\ \mathcal{S}_{k}^{\mathcal{F}}\left(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota)}\right),$ (32a)
@@
$\theta_{n}^{(\iota+1),k}\triangleq 2\pi-\angle{\boldsymbol{m}}^{(\iota+1)}_{k}(n),n=1,\dots,N,$ (33)
then we find $\boldsymbol{\theta}^{(\iota+1)}$ that maximize the $\mbox{SR}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta})$ and satisfies the UMC constraint (15c),
$\boldsymbol{\theta}^{(\iota+1)}\triangleq\arg\underset{\boldsymbol{\theta}\in\left\{\boldsymbol{\theta}^{(\iota+1),k},k=1,\dots,\mathcal{K}\right\}}{\max}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}).$ (34)
Algorithm 1 Proposed Iterative Algorithm for Solving Problem $(\mathcal{P}1)$
1:Initialize: $(\boldsymbol{w}^{(1)}=\boldsymbol{w}^{*},\boldsymbol{\theta}^{(1)}=\boldsymbol{\theta}^{*})$, convergence tolerance $\epsilon_{t}>0$, and Set $\iota=1$.
2:Repeat
3:Update $\boldsymbol{w}^{(\iota+1)}$ by (23), and $\boldsymbol{\theta}^{(\iota+1)}$ by (34);
4:if $\frac{|\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota+1)},\boldsymbol{\theta}^{(\iota+1)})-\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})|}{\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(\iota)},\boldsymbol{\theta}^{(\iota)})}\leq\epsilon_{t}$.
5:Then $\boldsymbol{\theta}^{(\iota)}\leftarrow\boldsymbol{\theta}^{(\iota+1)}$, $\boldsymbol{w}^{(\iota)}\leftarrow\boldsymbol{w}^{(\iota+1)}$ and terminate.
6:Otherwise $\iota\leftarrow\iota+1$ and continue.
7:Output $(\boldsymbol{w}^{*}=\boldsymbol{w}^{(\iota)}_{k},\boldsymbol{\theta}^{*}=\boldsymbol{\theta}^{(\iota)})$.
@@
$\mathbf{X}_{k}\triangleq\mathbf{\Phi}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}_{k}\boldsymbol{w}_{k}^{(\iota),H}\boldsymbol{G}_{{}_{\text{AR}}}^{H}\mathbf{\Phi}^{(\iota),H}+\mathbf{\Phi}^{(\iota)}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}_{k}^{(\iota)}\boldsymbol{w}_{k}^{H}\boldsymbol{G}_{{}_{\text{AR}}}^{H}\mathbf{\Phi}^{(\iota),H}-\mathbf{\Phi}^{(\iota)}\boldsymbol{G}_{{}_{\text{AR}}}\boldsymbol{w}_{k}^{(\iota)}\boldsymbol{w}_{k}^{(\iota),H}\boldsymbol{G}_{{}_{\text{AR}}}^{H}\mathbf{\Phi}^{(\iota),H},$ (40)
$\Delta\boldsymbol{u}_{k}\mathbf{X}_{k}\Delta\boldsymbol{u}_{k}^{H}+2\Re\{(\widehat{\boldsymbol{u}}_{k}^{H}\mathbf{X}_{k})\Delta\boldsymbol{u}_{k}\}+d_{k}\geq(2^{\varphi_{k}}-1)\beta_{k},\forall\|\Delta\boldsymbol{u}_{k}\|_{2}\leq\xi_{k},\forall k.$ (41)

### v3

@@ III-1 Sub-Problem for Optimizing the Beamforming Vectors
Finally, by substitution (22) in (16), and introducing an auxiliary variable $\mathcal{D}$ as the lower bound of the FBR-SR, we can recast problem ($\mathcal{P}1.1$) as:
$(\mathcal{P}1.3):$ $\max_{\Gamma,\boldsymbol{w}}\mathcal{D},$ (23a)
s.t. $\mathcal{D}\leq\widetilde{\mathcal{S}}_{k}^{\mathcal{F}}(\boldsymbol{w},\boldsymbol{\theta}^{(n)}),\leavevmode\nobreak\ \forall k,$ (23b)
[…]
@@
problem ($\mathcal{P}1.3$) is a Semidefinite Programming problem (SDP), with its linearized objective function in (22). This problem can be solved using standard solvers, e.g., the interior point method or the CVX toolbox [31].
@@ III-2 Sub-Problem for Optimizing the PREs
Likewise, given $\boldsymbol{w}$, we aim to find $\boldsymbol{\theta}^{(n+1)}$ such that, $\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(n+1)}_{k},\boldsymbol{\theta}^{(n+1)})\geq\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(n+1)}_{k},\boldsymbol{\theta}^{(n)})$. Similar to the previous section, the lower bound approximation of the user’s SR can be obtained using the inequality (71) in Appendix D. The user’s data rate can be expressed as in (24), where
${x}_{1,k}^{(n+1)}$ $\triangleq C_{k}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n)})-\gamma_{k}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n)})-\sigma_{k}{z}_{1,k}^{(n+1)},$
${z}_{1,k}^{(n+1)}$ $\triangleq{1}/{\rho_{k}^{(n+1)}}-{1}/{\upsilon_{{}_{k}}^{(n+1)}},$
$\boldsymbol{y}^{(n+1)}_{k,k}(m)$ $\triangleq\tilde{\boldsymbol{y}}^{(n+1)}_{k,k}(m)/{\rho_{k}^{(n+1)}},$
$\tilde{\boldsymbol{y}}^{(n+1)}_{k,k}(m)$ $\triangleq\left(\boldsymbol{w}^{(n+1)}_{k}\right)^{H}\boldsymbol{h}_{{}_{k}}^{H}(\boldsymbol{\theta}^{(n)})\boldsymbol{l}_{k}\boldsymbol{\Gamma}_{m}\boldsymbol{L}_{\text{AR}}\boldsymbol{w}^{(n+1)}_{k}$,
$\boldsymbol{\varphi}_{1,k}^{(n+1)}$ $\triangleq-{z}_{1,k}^{(n+1)}{\sum}_{j=1}^{\mathcal{K}}\boldsymbol{\varphi}_{k,j}^{(n+1)},$
$\boldsymbol{\varphi}_{k,j}^{(n+1)}$ $\triangleq\left(\boldsymbol{\Lambda}_{k,j}^{(n+1)}(m)\right)^{*}\boldsymbol{\Lambda}_{k,j}^{(n+1)}(\dot{m}),m\in M,\dot{m}\in M,$
$\boldsymbol{\Lambda}_{k,j}^{(n+1)}(m)$ $\triangleq\boldsymbol{l}_{k}\boldsymbol{\Gamma}_{m}\boldsymbol{L}_{{}_{\text{AR}}}\boldsymbol{w}^{(n+1)}_{j}$.
Algorithm 1 Proposed Iterative Algorithm for Solving $(\mathcal{P}1)$
1:Initialize: $(\boldsymbol{w}^{(1)}=\boldsymbol{w}^{*},\leavevmode\nobreak\ \boldsymbol{\theta}^{(1)}=\boldsymbol{\theta}^{*}$,  $\tau_{\max}$$\tau_{e_{\max}}$$t_{\max}$), convergence tolerance $\epsilon_{t}>0$, and Set $n=1$.
2:Repeat
3:Update $\boldsymbol{w}^{(n+1)}$ by (23), and $\boldsymbol{\theta}^{(n+1)}$ by (34);
4:if $\frac{|\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n+1)})-\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})|}{\underset{k\in\mathcal{K}}{\min}\mathcal{S}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(n)},\boldsymbol{\theta}^{(n)})}\leq\epsilon_{t}$.
5:Then $\boldsymbol{\theta}^{(n)}\leftarrow\boldsymbol{\theta}^{(n+1)}$, $\boldsymbol{w}^{(n)}\leftarrow\boldsymbol{w}^{(n+1)}$ and terminate.
6:Otherwise $n\leftarrow n+1$ and continue.
7:Output $(\boldsymbol{w}^{*}=\boldsymbol{w}^{(n)}_{k},\boldsymbol{\theta}^{*}=\boldsymbol{\theta}^{(n)})$.
@@
$x_{1,k_{e}}^{(n+1)}$ $\triangleq\ln\left(\rho_{e}^{(n+1)}\right)-\rho_{e}^{(n+1)}-\ln\left(\upsilon_{e}^{(n+1)}\right)+1,$
$z_{1,k_{e}}^{(n+1)}$ $\triangleq\left(-{\rho_{e}^{(n+1)}}/{1+\rho_{e}^{(n+1)}}-{1}/{\upsilon_{e}^{(n+1)}}\right)\sigma_{e},$
${\boldsymbol{y}}^{(n+1)}_{k_{e}}(m)$ $\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}{\boldsymbol{y}}^{(n+1)}_{j,k_{e}}(m),$
${\boldsymbol{y}}^{(n+1)}_{j,k_{e}}(m)$ $\triangleq\left(\boldsymbol{w}^{(n+1)}_{k}\right)^{H}\boldsymbol{h}_{e}^{H}(\boldsymbol{\theta}^{(n)})\boldsymbol{l}_{e}\boldsymbol{\Gamma}_{m}{\boldsymbol{L}}_{{}_{\text{{AR}}}}\boldsymbol{w}^{(n+1)}_{k}$,
$\boldsymbol{\varphi}_{1,k_{e}}^{(n+1)}$ $\triangleq-\dfrac{\rho_{e}^{(n+1)}{\sum}_{j=1}^{\mathcal{K}}{\boldsymbol{\varphi}}_{k_{e},j}^{(n+1)}}{(1+\rho_{e}^{(n+1)})}-\frac{{\sum}_{j=1}^{\mathcal{K}}{\boldsymbol{\varphi}}_{k_{e},j}^{(n+1)}}{(\sigma_{e}+\upsilon_{e}^{(n+1)})},$
$\boldsymbol{\varphi}_{k_{e},j}^{(n+1)}$ $\triangleq\left({\boldsymbol{\Lambda}}_{k_{e},j}^{(n+1)}(m)\right)^{*}\boldsymbol{\Lambda}_{k_{e},j}^{(n+1)}(\dot{m}),$
$\boldsymbol{\Lambda}_{k_{e},j}^{(n+1)}(m)$ $\triangleq\boldsymbol{l}_{e}\boldsymbol{\Gamma}_{m}{\boldsymbol{L}}_{{}_{\text{{AR}}}}\boldsymbol{w}^{(n+1)}_{j}$.
Next, we express the user-$k$’s and Eve dispersion factor. For $i\in\{k,e\}$, we can express the dispersion factors as in (26), where $x_{i,k}^{(n+1)}$ is expressed as in (27), $\boldsymbol{f}_{i,j}^{(n+1)}(m)$ is expressed as in (28), and
$\boldsymbol{f}_{i}^{(n+1)}$ $\triangleq{\sum}_{j=1,j\neq k}^{\mathcal{K}}\boldsymbol{f}_{i,j}^{(n+1)},$
$\boldsymbol{\varphi}_{2,i}^{(n+1)}$ $\triangleq{z}_{2,i}^{(n+1)}{\sum}_{j=1}^{\mathcal{K}}\boldsymbol{\varphi}_{i,j}^{(n+1)},$
${z}_{2,i}^{(n+1)}$ $\triangleq\dfrac{(\xi_{i}\rho^{(n+1)}_{i})}{\left(\upsilon^{(n+1)}_{i}\right)^{2}\sqrt{V_{i}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}^{(n)})}}.$
@@
$\widetilde{\mathcal{S}}_{k}^{\mathcal{F}}\left(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}\right)$ $\geq x_{k}^{(n+1)}+2{\sum}_{m=1}^{M}\Re\left\{\boldsymbol{y}^{(n+1)}_{k}(m)e^{j\theta_{m}}\right\},$ (29)
where $x_{k}^{(n+1)}$ is expressed as in (30), ${\boldsymbol{y}}^{(n+1)}_{k}(m)$ is expressed as in (31), and
$\boldsymbol{\varphi}_{k}^{(n+1)}$ $\triangleq\boldsymbol{\varphi}_{1,k}^{(n+1)}+\boldsymbol{\varphi}_{1,k_{e}}^{(n+1)}+\boldsymbol{\varphi}_{k_{e},j}^{(n+1)}+\boldsymbol{\varphi}_{2,k}^{(n+1)}+\boldsymbol{\varphi}_{2,k_{e}}^{(n+1)}.$
With (29), we formulate the following optimization problem to obtain $\boldsymbol{\theta}^{(n+1)}$,
@@
$\theta_{m}^{(n+1),k}\triangleq 2\pi-\angle{\boldsymbol{y}}^{(n+1)}_{k}(m),m=1,\dots,M,$ (33)
then we find $\boldsymbol{\theta}^{(n+1)}$ that maximize the $\mbox{SR}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta})$, and satisfies the UMC constraint (15c),
$\boldsymbol{\theta}^{(n+1)}\triangleq\arg\underset{\boldsymbol{\theta}\in\left\{\boldsymbol{\theta}^{(n+1),k},k=1,\dots,\mathcal{K}\right\}}{\max}\widetilde{\mathcal{S}}_{k}^{\mathcal{F}}(\boldsymbol{w}^{(n+1)},\boldsymbol{\theta}).$ (34)
@@
$\mathbf{A}_{k}\triangleq\mathbf{\Phi}\boldsymbol{L}_{\text{AR}}\boldsymbol{w}_{k}\boldsymbol{w}_{k}^{(n),H}\boldsymbol{L}_{\text{AR}}^{H}\mathbf{\Phi}^{(n),H}+\mathbf{\Phi}^{(n)}\boldsymbol{L}_{\text{AR}}\boldsymbol{w}_{k}^{(n)}\boldsymbol{w}_{k}^{H}\boldsymbol{L}_{\text{AR}}^{H}\mathbf{\Phi}^{(n),H}-\mathbf{\Phi}^{(n)}\boldsymbol{L}_{\text{AR}}\boldsymbol{w}_{k}^{(n)}\boldsymbol{w}_{k}^{(n),H}\boldsymbol{L}_{\text{AR}}^{H}\mathbf{\Phi}^{(n),H},$ (40)
$\Delta\boldsymbol{l}_{k}\mathbf{A}_{k}\Delta\boldsymbol{l}_{k}^{H}+2\Re\{(\widehat{\boldsymbol{l}}_{k}^{H}\mathbf{A}_{k})\Delta\boldsymbol{l}_{k}\}+d_{k}\geq(2^{\beta_{k}}-1)\alpha_{k},\forall\|\Delta\boldsymbol{l}_{k}\|_{2}\leq\Omega_{k},\forall k.$ (41)